In a study of cell phone usage and brain hemispheric dominance, an Internet survey was e-mailed to 6958 subjects randomly selThe P-value is (Round to three decimal places as needed.) Because the P-value is the significance level, the null hypothesis.
Answer #1

Solution :

Given that,

Po = 0.20

1 - Po = 0.80

n = 6958

x = 1286

Level of significance = \alpha = 0.01

Point estimate = sample proportion = \hat p = x / n = 0.185

This a left (One) tailed test.

The null and alternative hypothesis is,

F)

Ho: p = 0.20

Ha: p < 0.20

Test statistics

z = (\hat p - Po ) / \sqrt{} Po*(1-Po) / n

= ( 0.185 - 0.20) / \sqrt{} (0.20*0.80) / 6958

= -3.17

P-value = P(Z < z )

= P(Z < -3.17 )

= 0.001

Since, P-value < 0.01, We Reject H0.

Because the P-value is less than the significance level, We reject the null hypothesis. There is enough evidence to support the claim that the return rate is less than 20%.